# Chirality-induced spin selectivity based on orbital Edelstein effect in an analytically solvable model

### v1

@@ Abstract
Chirality-induced spin selectivity, a phenomenon wherein chiral structures selectively determine the spin polarization of electron currents flowing through the material, has garnered significant attention due to its potential applications in areas such as spintronics, enantioseparation, and catalysis. The underlying physical effect is the Edelstein effect that converts charge to angular momentum but the precise mechanism remains yet to be understood. Here, we introduce the minimal model for explaining the phenomenon based on the orbital Edelstein effect. We consider inter-site contributions to the current-induced orbital angular momentum and reveal the underlying mechanism by analytically calculating the Edelstein susceptibilities in a tight-binding and Boltzmann approach. While the orbital angular momentum is directly generated by the chirality of the crystal, the spin contribution of each spin-split band pair relies on spin-orbit coupling. Using tellurium as an example, we show that the orbital contribution surpasses the spin contribution by orders of magnitude.
@@
Chirality is a geometric property that affects a wide range of scientific fields, from chemistry and biology to physics and material engineering. A structure is chiral if it is distinguishable from its mirror image; it lacks mirror, inversion and roto-inversion symmetries [1]. The structural chirality of molecules like the DNA molecule or crystals like tellurium and selenium [cf. Fig. 1] can affect biochemical properties and even couples to physical observables like the angular momentum which gives rise to an intriguing phenomenon known as chirality-induced spin selectivity (CISS). The chirality of a structure determines the sign of the spin-polarization of electrons moving through it. CISS was first observed in chiral molecules [2]. More recently, this phenomenon has been observed in a broader range of systems, including hybrid organic-inorganic perovskites [3, 4] and chiral crystals like tellurium [5, 1, 6] or CrNb3S6 [7, 8].
@@
Despite extensive experimental observations, the underlying physical mechanisms governing CISS remain a topic of active debate. Theoretical explanations are mostly based on the charge-to-spin conversion effect named inverse spin-galvanic effect or Edelstein effect (EE) [9, 10]: Due to a broken inversion symmetry, the spin polarizations of electrons with opposite wave vectors differ which leads to the generation of a spin density once an electric field or currents are applied. The EE typically occurs at interfaces where the broken inversion symmetry gives rise to a Rashba-type spin-orbit coupling [11, 12, 13]. In a chiral material, however, the inversion symmetry is innately broken and a collinear effect can occur where the spins are parallel to the wave vector [5, 1, 6, 14, 15] and the theoretically calculated Edelstein susceptibilities indeed depend on the chirality of the structure. However, as was pointed out in Refs. [14, 15], they are an order of magnitude smaller than what was measured experimentally in tellurium [5] pointing towards the existence of important additional contributions.
@@
Recently, several phenomena related to angular momentum have been observed and understood based on the orbital degree of freedom. Orbital-polarized currents, generated by the orbital Hall effect [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] or the orbital EE [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] can lead to the generation or transport of orbital angular momentum when currents or electric fields are applied. It has been shown that the orbital EE is sensitive to the chirality of a system as well [33, 34, 38, 39, 42]. We refer to this phenomenon as chirality-induced orbital selectivity (CIOS). When a current flows through a chiral material, it becomes orbital polarized due to a shift of the Fermi surfaces, as visualized in Fig. 2. Once the current enters the electrodes, like Pt which has a considerable spin-orbit coupling (SOC), it is transformed into a spin-polarized current. CIOS is therefore detected effectively as CISS in most experiments, where the orbital contribution is simply scaled by the orbital-to-spin conversion efficiency of the electrode. This contribution occurs additionally to the genuine spin contribution to CISS and could account for the discrepancy of theoretically and experimentally determined CISS efficiencies.
@@
While this has been an important achievement in understanding CISS, so far, the precise mechanism of CIOS is not well understood. Since in most theoretical descriptions, orbital angular momentum is calculated via the hybridization of multiple atomic orbitals at the same atom (atomic-center approximation), the efficiency of CIOS is underestimated and can hardly be understood intuitively because the genuine chiral mechanism is not captured: The orbital angular momentum is mostly generated by the circular motion of electrons among multiple atoms forming a helix, like in tellurium. This generates an orbital angular momentum that is always along the current direction for one helix chirality and always opposite for the other (cf. arrows in Fig. 1). Currently, the literature on the orbital Edelstein effect considering such inter-site contributions is scarce. Refs. [33, 34] use the modern formulation of orbital magnetization  [45, 46, 47, 48, 49, 50] and even make the analogy to a classical solenoid. However, due to the Haldane-like [51] nature of the model [33, 34], requiring multiple independent parameters, the Edelstein susceptibility cannot be calculated analytically. Furthermore, the interpretation based on Weyl points in the three-dimensional Brillouin zone dilutes the intuitive mechanism based on the chirality in real space.
@@
In this paper, we introduce the minimal model for understanding CIOS and CISS. The unit cell of the considered helix consists of only three $s$ orbitals and extends periodically along one dimension. We consider the modern formulation of orbital magnetization for the calculation of the orbital angular momentum to account for inter-site contributions that are crucial for describing the effect in chiral materials properly. Since we can describe this system analytically, we are able to reveal the chiral mechanism of CIOS and CISS based on an orbital EE that is orders of magnitude larger than the spin EE.
@@
Figure 1: Chiral helices as in tellurium. (a) Right-handed helix structure characterized by $c>0$. (b) Left-handed structure characterized by $c<0$. The black arrow symbolizes the translational motion of electrons along $z$. It is constrained to a circular motion (red and blue) which generates opposite orbital angular momentum $L_{z}$ for the two opposite crystal chiralities.
Minimal model.
@@
The chiral structure is considered periodic along the $z$ direction with a period of $c$. The sign of $c$ controls the chirality of the helix: right-handed for $c>0$ and left-handed for $c<0$, cf. Fig. 1. We also consider periodicity in the $xy$ plane but for simplicity the helices are not coupled by hoppings. This allows us to describe the chiral structure in an effectively one-dimensional Brillouin zone with $k_{z}$ the only wave vector component that exhibits dispersion.
@@
This band structure [Fig. 3(a)] is equivalent to the single band $-2t\cos\left(k\,\frac{c}{3}\right)$ characterizing a one-dimensional chain of atoms with lattice constant $c/3$ along $z$ which has been back-folded into the smaller Brillouin zone accounting for the $3$ atoms in the unit cell. The eigenvectors $\lvert n\rangle$ along $k_{z}$ are constant and read
@@
This results in a chirality-dependent orbital angular momentum (opposite for $c>0$ versus $c<0$)
@@
that is proportional to the group velocity $v_{n,z}(k)=\frac{1}{\hbar}\frac{\partial E_{n}}{\partial k_{z}}$. The fact that the orbital angular momentum is antisymmetric with $k$, while the band structure is symmetric [cf. also Fig. 3(a)], gives rise to an orbital Edelstein effect, as visualized in Fig. 2.
The collinear orbital Edelstein susceptibility $\chi_{z}^{L_{z}}$ relates the applied electric field $\boldsymbol{E}=E_{z}\boldsymbol{e}_{z}$ and the generated non-equilibrium orbital magnetic moment per unit cell $m_{L_{z}}^{\mathrm{uc}}$ via $m_{L_{z}}^{\mathrm{uc}}=\chi_{z}^{L_{z}}E_{z}$. We can calculate $\chi_{z}^{L_{z}}$ easily using the Boltzmann transport theory when we assume a constant relaxation time $\tau$
@@
So far, we wanted to make the discussion as general as possible. Now we consider the case of tellurium’s $s$ bands that exhibit an almost perfect resemblance of the band structure we have calculated. We use $a=2.133\,$Å, $t=1.249\,\mathrm{eV}$ [53] and $g_{L}=1$ which gives a maximum value for the orbital angular momentum $L_{0}=0.43\,\hbar$. For the orbital Edelstein susceptibility we additionally use $c=\pm 6.018\,$Å  and assume a typical relaxation time $\tau=1\,\mathrm{ps}$ which results in an extremum of $\chi_{0}=\pm 125.0\times 10^{-9}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. The sign depends on the crystal chirality which means the system exhibits an extremely strong yet purely geometrical CIOS that is converted into CISS once measured. $\chi_{z}^{L_{z}}$ is orders of magnitude larger than the spin Edelstein susceptibility $\chi_{z}^{S_{z}}$ calculated for tellurium in Ref. [6]. In our model $\chi_{z}^{S_{z}}=0$, because spin is not yet considered.
@@
We note that we have restricted our model to one dimension, so that we can solve it analytically. In reality, the individual helices are coupled by van der Waals interactions and dispersion along $k_{x}$ and $k_{y}$ emerges. However, especially near the center of the energy range, the Fermi surfaces are sheets that extend throughout the whole $k_{x}k_{y}$ plane of the three-dimensional Brillouin zone, similar to the schematics shown in Fig. 2(b). In Figs. S1, S2 in the Supplemental Material [54] we show that the full three-dimensional system is here described well by the quasi-one-dimensional analytical model.
Effect of spin-orbit coupling.
After having established the main result of the paper, next, we present how including SOC influences the orbital Edelstein effect and may cause a spin Edelstein effect as well; two effects relying on distinct mechanisms. First, we make the Hamiltonian spin dependent. Without SOC, all bands are now doubly degenerate but the orbital angular momentum always has the same value for the two uncoupled spin-subsystems. Therefore, the orbital Edelstein susceptibility $\chi_{z}^{L_{z}}(E)$ is simply increased by a factor of 2 to a maximum of $250.0\times 10^{-9}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. The spin Edelstein susceptibility $\chi_{z}^{S_{z}}(E)=0$ due to Kramer’s degeneracy.
SOC is included akin to the Kane-Mele model [55], that can be written in second quantization
$i\lambda\sum_{\braket{i,j}}\nu_{ij}c_{i}^{\dagger}\sigma_{z}c_{j}.$ (8)
Here $c_{i}^{\dagger}$ and $c_{j}$ are the creation and annihilation operators at sites $i,j$. $\nu_{ij}=\pm 1$ accounts for the hopping path along the helix: $+1$ for a counter-clockwise rotation and $-1$ for a clockwise rotation, so $+\mathrm{sign}(c)$ for hopping along $+z$ and $-\mathrm{sign}(c)$ for hopping along $-z$. The resulting band structure [Figs. 4(a,b)] is
Note the sign change of the SOC term for $c>0$ and $c<0$: The sine function comes from adding the two hopping terms along ±$\pm c\boldsymbol{e}_{z}$ where the upwards hopping path along the helix is rotating oppositely to the downwards hopping path: $i[\exp(ick_{z})-\exp(-ick_{z})]$. Furthermore, it depends on the spin-orientation. Close to the $\Gamma$ point the SOC-induced term is $\mp 2\lambda k_{z}c/3$, a term that has been derived in Ref. [56] as well.
@@
The Hamiltonian is separated into spin-up and spin-down blocks meaning that all bands are ±$\pm 100\,\%$ spin polarized $S_{n,z}^{\uparrow\downarrow}(k)=\pm\frac{\hbar}{2}$ [color in Fig. 4(b)]. This makes the system under consideration similar to a Dresselhaus system [57], where the structural chirality breaks the inversion symmetry, but in one dimension. The spin is (anti-) parallel to $k$; a feature well known from the Fermi surface of tellurium so the expected spin Edelstein effect is collinear [5, 1, 6, 14, 15] meaning that electric field and generated magnetic moment are collinear.
Conversely, the orbital angular momenta of the two bands of each spin-split band pair are almost the same [Fig. 4(a)]. They are affected in a similar way as the band structure
Even if the SOC were $5\%$ of the hopping amplitude $\lambda=0.05t$, the maximum of the orbital Edelstein susceptibility changes by only $0.125\%$. This highlights that the orbital Edelstein effect is not governed by SOC but purely by the chiral structure of the helices.
The contributions of the inner and outer bands to the spin Edelstein susceptibility cancel, $\chi_{z}^{S_{z}}=0$. In Fig. S3 in the Supplemental Material [54], we show that once a spin EE emerges, e. g. by considering a $k$-dependent relaxation time, it is proportional to spin-orbit coupling and orders of magnitude smaller than the orbital EE.
Discussion.
We have analytically shown that the current-induced selectivity of spin and orbital angular momentum in chiral materials like selenium and tellurium follow fundamentally different mechanisms. The orbital contributions from bands with opposite spin directions add up and the effect occurs even without SOC. The spin EE of both bands (almost) cancels and scales with SOC making it a less efficient mechanism. As mentioned in the introduction, in an experiment, the orbital-polarized currents will be converted into spin-polarized currents upon passing through the detecting electrode so the CIOS will effectively be detected as CISS. Our work fills the gap between large experimentally observed signals and the Edelstein signals that were calculated before based on only the spin EE or on the orbital EE effect in atomic-center approximation.
Effective single-band description.

### v2

# Chirality-induced orbital Edelstein effect in an analytically solvable model

@@ Abstract
Chirality-induced spin selectivity (CISS), a phenomenon wherein chiral structures selectively determine the spin polarization of electron currents flowing through the material, has garnered significant attention due to its potential applications in areas such as spintronics, enantioseparation, and catalysis. The underlying physical effect is the Edelstein effect that converts charge to angular momentum. Besides a spin contribution there exists a contribution based on the orbital angular momentum but the precise mechanism for its generation remains yet to be understood. Here, we introduce the minimal model for explaining the phenomenon based on the orbital Edelstein effect. We consider non-local inter-site contributions to the current-induced orbital angular momentum and reveal the underlying mechanism by analytically calculating the Edelstein susceptibilities in a tight-binding and Boltzmann approach. While the orbital angular momentum is directly generated by the chirality of the crystal, the spin contribution of each spin-split band pair relies on spin-orbit coupling. Using tellurium as an example, we show that the orbital contribution surpasses the spin contribution by orders of magnitude.
I Introduction
@@
Chirality is a geometric property that affects a wide range of scientific fields, from chemistry and biology to physics and material engineering. A structure is chiral if it is distinguishable from its mirror image; it lacks mirror, inversion and roto-inversion symmetries [1]. The structural chirality of molecules like the DNA molecule or crystals like tellurium and selenium [cf. Fig. 1] can affect biochemical properties and even couples to physical observables like the angular momentum which gives rise to intriguing phenomena like electrical magnetochiral anisotropy [2, 3, 4, 5] and chirality-induced spin selectivity (CISS) [6, 7, 8, 1, 9, 5, 10, 11, 12, 13, 14]. The chirality of a structure determines the sign of the spin-polarization of electrons moving through it. CISS was first observed in chiral molecules [6]. More recently, this phenomenon has been observed in a broader range of systems, including hybrid organic-inorganic perovskites [7, 8] and chiral crystals like tellurium [15, 1, 9] or CrNb3S6 [5, 10].
@@
Despite extensive experimental observations, the underlying physical mechanisms governing CISS remain a topic of active debate. Some theoretical explanations are based on the charge-to-spin conversion effect named inverse spin-galvanic effect or Edelstein effect (EE) [16, 17]: Due to a broken inversion symmetry, the spin polarizations of electrons with opposite wave vectors differ which leads to the generation of a spin density once an electric field or currents are applied. The EE typically occurs at interfaces where the broken inversion symmetry gives rise to a Rashba-type spin-orbit coupling [18, 19, 20]. In a chiral material, however, the inversion symmetry is innately broken and a collinear effect can occur where the spins are parallel to the wave vector [15, 1, 9, 21, 22] and the theoretically calculated Edelstein susceptibilities indeed depend on the chirality of the structure. However, as was pointed out in Refs. [21, 22], they are an order of magnitude smaller than what was measured experimentally in tellurium [15] pointing towards the existence of important additional contributions.
@@
Recently, several phenomena related to angular momentum have been observed and understood based on the orbital degree of freedom. Orbital-polarized currents, generated by the orbital Hall effect [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] or the orbital EE [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51] can lead to the generation or transport of orbital angular momentum when currents or electric fields are applied. It has been shown that the orbital EE is sensitive to the chirality of a system as well [40, 41, 45, 46, 49]. We refer to this phenomenon as chirality-induced orbital selectivity (CIOS). When a current flows through a chiral material, an orbital magnetic moment is generated due to a shift of the Fermi surfaces, as visualized in Fig. 2. If the chiral material is interfaced with another material, an orbital current may flow. This material may serve as an electrode, which typically has a considerable spin-orbit coupling (SOC), for example Pt. This leads to the conversion of the orbital current into a spin current. The orbital Edelstein effect may then be detected as CISS and not as CIOS, where the orbital contribution is scaled by the orbital-to-spin conversion efficiency of the electrode. This contribution occurs additionally to any genuine spin contribution to CISS and could therefore account for the discrepancy of theoretically and experimentally determined CISS efficiencies. Therefore, understanding the mechanism behind the orbital Edelstein effect in chiral materials is of crucial importance.
@@
In most theoretical descriptions, orbital angular momentum is calculated via the hybridization of multiple atomic orbitals at the same atom (atomic-center approximation), the efficiency of CIOS is underestimated and can hardly be understood intuitively because the genuine chiral mechanism is not captured: The orbital angular momentum is mostly generated by the circular motion of electrons among multiple atoms forming a helix, like in tellurium. This generates an orbital angular momentum that is always along the current direction for one helix chirality and always opposite for the other (cf. arrows in Fig. 1). Even in a semi-classical analysis, we can understand why considering inter-side hybridization is crucial: Electrons moving along a helix generate a current and orbital angular momentum along the same direction. The orbital-polarized current is independent of the motion direction but changes sign with the chirality of the helix because the angular velocity is proportional to the velocity along the helix and the chirality (cf. Fig. 1). Here it becomes apparent that the orbital angular momentum is generated due to the circular motion among multiple atoms forming the helix and is not introduced by hybridization of orbitals of one atom. Currently, the literature on the orbital Edelstein effect considering such inter-site contributions is scarce. Refs. [40, 41] use the modern formulation of orbital magnetization  [52, 53, 54, 55, 56, 57] and even make the analogy to a classical solenoid. However, due to the Haldane-like [58] nature of the model [40, 41], requiring multiple independent parameters, the Edelstein susceptibility cannot be calculated analytically. Furthermore, the interpretation based on Weyl points in the three-dimensional Brillouin zone dilutes the intuitive mechanism based on the chirality in real space.
@@
In this paper, we introduce the minimal model for understanding the Edelstein effect in chiral materials. The unit cell of the considered helix consists of only three $s$ orbitals and extends periodically along one dimension. We consider the modern formulation of orbital magnetization for the calculation of the orbital angular momentum to account for inter-site contributions that are crucial for describing the effect in chiral materials properly. Since we can describe this system analytically, we are able to reveal the chiral mechanism of CIOS and CISS based on an orbital EE that is orders of magnitude larger than the spin EE.
@@
Figure 1: Chiral helices. (a) Right-handed helix structure characterized by $c>0$. (b) Left-handed structure characterized by $c<0$. The black arrow symbolizes the translational motion of electrons along $z$. It is constrained to a circular motion (red and blue) which generates opposite orbital angular momentum $L_{z}$ (orange and green) for the two opposite crystal chiralities.
This paper is structured as follows. First, we introduce the model and analytically calculate the band structure, orbital angular momentum and the orbital Edelstein susceptibility in Sec. II. Afterwards, in Sec. III, we discuss the consequences of two important effects present in real materials: dispersion perpendicular to the helix direction and spin-orbit coupling. In Sec. IV, we show that the analytical model can be even further simplified by making use of a constraint between the angular velocity and the velocity along the helix, as in a classical analysis. We discuss our results and conclude in Sec. V.
II Orbital Edelstein effect in analytically solvable model
@@
The chiral structure is considered periodic along the $z$ direction with a period of $c$. The sign of $c$ controls the chirality of the helix: right-handed for $c>0$ and left-handed for $c<0$, cf. Fig. 1. We also consider periodicity in the $xy$ plane but for simplicity the helices are not coupled by hoppings. This allows us to describe the chiral structure in an effectively one-dimensional Brillouin zone with $k_{z}$ the only wave vector component that exhibits dispersion. The $x$ and $y$ coordinates of the atoms are important to account for the chirality and for the definition of the orbital angular momentum.
@@ II.1 Hamiltonian and band structure
This band structure [Fig. 3(a)] is equivalent to the single band $-2t\cos\left(k\,\frac{c}{3}\right)$ characterizing a one-dimensional chain of atoms with lattice constant $c/3$ along $z$ which has been back-folded into the smaller Brillouin zone accounting for the 333 atoms in the unit cell. Note that the band structure can be very easily calculated by diagonalizing the Hamiltonian for $k_{x}=k_{y}=0$. The eigenvectors $\lvert n\rangle$ along $k_{z}$ are constant and read
@@ II.2 Orbital angular momentum
The derivatives of the Hamiltonian $\frac{\partial H}{\partial k_{x}}$, $\frac{\partial H}{\partial k_{y}}$ account for the chirality of the system. This results in a chirality-dependent orbital angular momentum (opposite for $c>0$ versus $c<0$)
@@
that is proportional to the group velocity $v_{n,z}(k)=\frac{1}{\hbar}\frac{\partial E_{n}}{\partial k_{z}}$.
II.3 Orbital Edelstein effect
The fact that the orbital angular momentum is antisymmetric with $k$, while the band structure is symmetric [cf. also Fig. 3(a)], gives rise to an orbital Edelstein effect, as visualized in Fig. 2. The collinear orbital Edelstein susceptibility $\chi_{z}^{L_{z}}$ relates the applied electric field $\boldsymbol{E}=E_{z}\boldsymbol{e}_{z}$ and the generated non-equilibrium orbital magnetic moment per unit cell $m_{L_{z}}^{\mathrm{uc}}$ via $m_{L_{z}}^{\mathrm{uc}}=\chi_{z}^{L_{z}}E_{z}$. We can calculate $\chi_{z}^{L_{z}}$ easily using the Boltzmann transport theory when we assume a constant relaxation time $\tau$
@@ II.3 Orbital Edelstein effect
So far, we wanted to make the discussion as general as possible. Now we consider the case of tellurium’s $s$ bands that exhibit an almost perfect resemblance of the band structure we have calculated. We use $a=2.133\,$Å, $t=1.249\,\mathrm{eV}$ [60] and $g_{L}=1$ which gives a maximum value for the orbital angular momentum $L_{0}=0.43\,\hbar$. For the orbital Edelstein susceptibility we additionally use $c=\pm 6.018\,$Å  and assume a relaxation time $\tau=1\,\mathrm{ps}$ which results in an extremum of $\chi_{0}=\pm 125.0\times 10^{-9}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. Note that the magnitude of this signal scales with the relaxation time which is a parameter in our calculations. The sign depends on the crystal chirality which means the system exhibits an extremely strong yet purely geometrical orbital Edelstein effect. In our model $\chi_{z}^{S_{z}}=0$, because spin is not yet considered.
@@
III Additional effects present in real materials
After having established the main result of the paper, we test its validity by comparison with a three-dimensional model (cf. Sec. III.1). Afterwards, we present how including SOC influences the orbital Edelstein effect and may cause a spin Edelstein effect as well; two effects relying on distinct mechanisms (cf. Sec. III.2).
III.1 Hopping between the helices
We note that we have restricted our model to quasi one dimension, so that we can solve it analytically. In reality, the individual helices are coupled by van der Waals interactions and dispersion along $k_{x}$ and $k_{y}$ emerges. However, especially near the center of the energy range, the Fermi surfaces are sheets that extend throughout the whole $k_{x}k_{y}$ plane of the three-dimensional Brillouin zone, similar to the schematics shown in Fig. 2(b). In Appendix A, we show that the full three-dimensional system is here described well by the quasi-one-dimensional analytical model.
III.2 Influence of spin-orbit coupling
The following discussion is again for the one-dimensional model discussed in Sec. II. To include the effect of SOC, first, we make the Hamiltonian spin dependent. Without SOC, all bands are now doubly degenerate but the orbital angular momentum always has the same value for the two uncoupled spin-subsystems. Therefore, the orbital Edelstein susceptibility $\chi_{z}^{L_{z}}(E)$ is simply increased by a factor of 2 to a maximum of $250.0\times 10^{-9}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. The spin Edelstein susceptibility $\chi_{z}^{S_{z}}(E)=0$ due to Kramer’s degeneracy.
@@ III.2 Influence of spin-orbit coupling
SOC is included via a spin- and chirality-dependent hopping term along the helix direction. We introduce the term in Appendix B and explain that it corresponds to a reduced version of the Kane-Mele SOC term [61]. The resulting band structure exhibits spin-split bands in which the energy is slightly shifted along the ±$\pm k_{z}$ direction. Likewise, the orbital angular momentum is also only slightly affected by this shift along ±$\pm k_{z}$ [cf. Fig. 4(a)]. The analytical calculations are presented in Appendix B.
[…]
IV Effective single-band description